# Catch Me If You Can: Finding the Source of Infections in Temporal Networks

### v1

@@ 1 Introduction
Information, diseases, pollutants—all these things spread through the networks that define how everything from people and servers to our streets and sewers are connected. The field of spreading processes has received substantial attention from researchers as diverse as physicians, physicists, sociologists, and computer scientists. One of the central questions that naturally arises is to find the source of some spreading process given only limited information (e.g., from a small set of disease infection times or from a number of sensors in a sewer network) sir-source-detection; shah2011rumors; sensor-placement; paluch_fast_2018. The current mainstream models in this area, such as the independent cascade model independent-cascade, or the susceptible-infected-resistant (SIR) model Hethcote1989; a-contribution; shah2011rumors incorporate the inherent temporal behavior of the spreading process, that is, infections themselves take time such that a node first gets infected and only then is able to infect other nodes. However, they fail to capture temporal dependencies in the underlying network itself. In many real-world networks the connections change over time (e.g., social networks in which diseases or information spread are highly temporal as the link between two people is only able to transmit information or diseases at specific points in time) and these changes heavily impact the behavior of spreading processes. For this reason, the study of infections in temporal graphs has been initiated. In 2015, Gayraud et al. gayraud-evolving-social-networks, generalized the independent model from static to temporal graphs. Since then, other papers have studied problems related to spreading processes in temporal networks influencers, but, to the best of our knowledge, no one has conducted research on the problem of source detection (SD). We provide the first rigorous definition of SD in temporal graphs and give both algorithms and lower bounds for this problem in a variety of settings.
@@ 1.2 Related Work
The SD problem has a rich literature and has even been studied on the SIR model we employ. In 2011, shah2011rumors study the SD problem under the SIR model. In 2016, sir-source-detection extend their work. Other notable algorithmic contributions include the PVTA algorithm proposed by pinto2012locating, which focuses on estimating the source of an infection from observations on a sparse set of nodes. This model has also recently been extended to SD pmlr-v216-berenbrink23a. In 2011, shah2011rumors study the SD problem under the SI model. sir-source-detection extend this work to the SIR model. gayraud-evolving-social-networks, extend the study of influence maximization on both the independent cascade and linear threshold models to temporal graphs (which were popularized by kempe2000connectivity). influencers study a number of variations of the influence maximization problem on temporal graphs under the SIS model (which is closely related to SIR). color=pink]@Supervisors: Discuss temporal graphs themselves in the related work section?
@@ 2 Preliminaries
We write $2\operatorname{\mathbb{N}}$ for the set of even natural numbers and $12\operatorname{\mathbb{N}}+$ 1 for the set of odd natural numbers. Two real-valued functions $f,g\colon\operatorname{\mathbb{N}}\to\operatorname{\mathbb{R}}$ are asymptotically equivalent if $(f(n)/g(n))\xrightarrow{n\to\infty}1$. We use the standard definition of treewidth as presented in cygan2015parameterized.
@@ 3 The Source Detection Game
color=pink]Fix that the Discoverer learns the infection source
@@ 4 Randomized Source Detection
Randomizing the behavior of the Discoverer enables us to evade the worst cases and yield good results with high probability.color=pink]@Supervisors: Should I include a very bad lower bound on worst-case performance to justify this? — Yes I think that would be very helpful. In fact, in this section, we even show that there is a randomized Discoverer that wins the SD game with consistent source behavior on unknown static graphs within $O(n\sqrt{n})$ infections and with constant success probability.
@@
In each round, the strategy picks a node to watch uniformly at random. It terminates when it has observed a node which gets infected while being watched. Let $T$ be the random variable that takes the number of the round in which this happens. Let $a_{1},\dots$ be the nodes infected in the respective rounds. Then the expected number of tolerated infections is $\mathbb{E}\left[\sum_{i=1}^{T}a_{i}\right]=\sum_{i=1}^{\infty}a_{i}\cdot\mathbb{P}\left[i\leq T\right]=\sum_{i=1}^{\infty}a_{i}\cdot\prod_{j=1}^{i-1}\left(\frac{n-a_{j}}{n}\right)$, where the equalities follow from the alternative definition of the expectation and since we require at least $i$ rounds iff the first $i-1$, rounds are unsuccessful. Now, substitute $p_{i}=a_{i}/n$ for all $i$. Then $p_{i}$ is the probability of finishing in round $i$ assuming it is reached. We may also pull $n$ outside the sum to view the number of tolerated infections in each round as a multiple of $n$. This leaves us with $n\sum_{i=1}^{\infty}p_{i}\cdot\prod_{j=1}^{i-1}(1-p_{j})$. Now group the rounds of our strategy and thus the $p_{i}$ by taking consecutive elements until their sum is at least $1/2$ and then start the next group. Formally, let $\{k_{j}\}_{j\in\operatorname{\mathbb{N}}}$ be the (possibly finite) sequence such that $k_{j}$ is the smallest integer such that $\sum_{i=k_{j}}^{k_{j+1}-1}p_{i}\geq 1/2$. Then, by union bound, the probability of finishing within a given group, assuming we reach it, is at least $1/2$. Also, we have that $\sum_{i=k_{j}}^{k_{j+1}-1}p_{i}\leq 3/2$ since all but the last elements together are less than $1/2$ and the last element is at most 1. This is a bound on the tolerated infections (as a multiple of $n$). Therefore, we can bound the above expectation from above by the geometric process that tolerates $23n/$ 2 infections to perform the rounds in a given group and has probability $1/2$ of finishing within that block. Thus, we can bound above by
[…]
@@
Let $10<\varepsilon<$ 1 be arbitrarily small. For any algorithm solving the SD problem on trees with consistent source behavior, unknown static graph and watching a single node with success probability $p>0$, there is an instance for every number of nodes $n$ such that the price of detection under that algorithm is at least $(1/2)\cdot\sqrt{\ln\left((1-p/(2+\varepsilon))^{-2}\right)}n\sqrt{n}$.
This also gives us a bound on the trade-off between the success probability and the price of detection. For a more intuitive bound, obverse that by simple analytic tools we can deduce that even a price of detection of $pn\sqrt{n}$ is unachievable for any algorithm with success probability $p<0.8724$. color=green]@Niko, recheck: I am confused about what you really want to say here. Doesn’t $\sqrt{\ln\left(\frac{1}{(1-p/(2+\varepsilon))^{2}}\right)}>p$ for $p\in(0,1)$ give you that reducing $p$ gives you a better than proportional payoff? Or not even that. {proofsketch} For $n\in\operatorname{\mathbb{N}}^{+}$, we consider the path on $n$ nodes labeled $\operatorname{\lambda}(i,i+1)\coloneqq i$ for $i\in[n-1]$. We assume that there is an algorithm $A$ that wins the game on these graphs within $c\sqrt{n}$ rounds with probability at least $p$. From $A$ we construct algorithm $B$ with strictly stronger capabilities: in each round, $B$ watches the same nodes as $A$ and also all nodes that $A$ has observed to become infected in previous rounds. We argue that for $A$ to win the game, $B$ has to win the game, which happens if $A$ guesses a node which is close to the source. We bound the probability of that event to obtain the result.
@@
First, we describe an infinite family of graphs, then assume there in an algorithm winning the game on all of these instances color=green]@Niko, recheck: what does it mean to win a game on a family of graphs? in $o(n\sqrt{n})$ infections with probability at least $p$. We then derive an algorithm that must solve the problem strictly faster and show that, for all small enough $c$ (which we will give later), there are problem instances where this improved algorithm has probability less than $p$ to require less than $c\sqrt{n}$ rounds.
First of all, let $n\in\operatorname{\mathbb{N}}$. Then set $P_{n}$ the path with $n$ nodes and let the time labeling be $\operatorname{\lambda}(i,i+1)=i$ for $i\in[n-1]$. We shuffle the node indices uniformly at random before providing them to the Discoverer so that they do not reveal the graph For simplicity we will refer to the ordered indices in the rest of the proof. Now let $A$ be any algorithm that solves the SD problem on the set $\{P_{n}\}_{n\in\operatorname{\mathbb{N}}}$. w.l.o.g., we may assume that there is a round in which $A$ watches the root node (if there is no such round, we can modify $A$ such that in a single extra round before submitting its answer, $A$ watches the node it will claim is the source).color=green]@Niko, recheck: Maybe we should mention that we shuffle the nodes so that the indices of them do not give anything away about the position? That is somewhat obvious but it might help to mention it explicitly.
@@
By assumption, $A$ requires less than $c\sqrt{n}$ rounds (with probability at least $p$ and for large enough $n$). We may also assume $A$ spends at least one round watching the source once it has found it. Thus, in order for $B$ to have picked the source at the end, $A$ must have picked a node with distance at most $c\sqrt{n}$ from the source at some time color=green]@Niko, recheck: That sentence is confusing. A must have actually picked the source itself, so where does the distance come from?. In order for that to happen, $A$ must have picked a node about which it had no information and which has distance at most $c\sqrt{n}$ from the source (by counting the rounds needed to trace an infected node to the source).
Since the nodes about which $A$ has no information were shuffled uniformly at random by the adversary, we can assume that $A$ selects them uniformly at random without replacement. Now, since $22+\varepsilon>$ 2, the probability of picking a node within distance $c\sqrt{n}$ to the source is smaller than
@@
If we now pick $c<(1/2)\cdot\sqrt{\ln\left(\frac{1}{(1-p/(2+\varepsilon))^{2}}\right)}$ and $n$ large enough such that $c\sqrt{n}\geq 1$, we have that this success probability is less than p. This follows since, by choice, $-2c^{2}>\ln(1-p/(2+\varepsilon))$ and thus $e^{-2c^{2}}=\left(e^{-2c}\right)^{c}\geq 1-p/(2+\varepsilon)$. As $\left(1-\frac{(2+\varepsilon)c}{\sqrt{n}}\right)^{\sqrt{n}}$ approaches $e^{-(2+\varepsilon)c}$ and $\frac{n-(2+\varepsilon)c\sqrt{n}}{n}=1-\frac{(2+\varepsilon)c}{\sqrt{n}}$, we have the desired result $1-\left(\frac{n-(2+\varepsilon)c\sqrt{n}}{n}\right)^{c\sqrt{n}}<p$ for large enough $n$. This contradicts the assumption that $A$ had success probability at least $p$, proving that no such algorithm may exist.
@@ 5 When Knowing the Static Graph Helps
The main idea of this reduction is to use the algorithm $A$ as a subroutine by reporting (${V\choose 2}$ as the edge set to it.
@@
Figure 3: The construction from the proof of Section 5. Discoverer for known graphs $A$ is used on the completed graph on the right to win the game in the setting of unknown graphs.
color=green]@Niko, recheck: It should be $T_{max}^{\prime}$ right?
@@
This result is achieved by Algorithm 2. It crucially depends on the existence of small separators. Specifically, in a static graph $G$ with treewidth $\mathrm{tw}{}$, there is a $1/2$-balanced separator $S$ (i.e., every connected component in $G-S$ has size at most $1/2\left|V(G)\right|$) of size at most $\mathrm{tw}+1$ cygan2015parameterized. See Figure 4 for an illustration. This result also holds if the nodes are weighted, and we ask for a separator to ensure each component has at most half of the weight. Therefore, if we assign either 00 or $1$ as weights, we may pick which nodes we want to evenly distribute on the sides of the separator.
@@ 7 Dynamic Infection Behavior
Let $n\in 2\operatorname{\mathbb{N}}+1$. We construct $G$ as follows: take $(n-1)/2$ edge-disjoint Hamiltonian paths $P_{1},\dots,P_{(n-1)/2}$ on the $K_{n-1}$ (which exist by axiotis_approx), add a node $s$ and connect $s$ to the first node on each of the Hamiltonian paths. From now on, we interpret each $P_{i}$ as starting at $s$. Order the nodes and the Hamiltonian paths uniformly at random. Note, these cover all edges in $\mathcal{G}$. Write $p\colon E(\mathcal{G})\to[n/2]$ be the function that maps an edge to the Hamiltonian path it belongs to and write $q\colon E(\mathcal{G})\to[n]$ for the function that assigns an edge its index on that path. Then let the Adversary act according to $\operatorname{\lambda}\coloneqq e\mapsto p(e)n+q(e),$ and pick $\delta=1$. See Figure 5 for an illustration of this construction. In each round $i$, the Adversary lets the source infect the whole graph via $P_{i\operatorname{mod}(n-1)/2}$. For the sake of this proof, we may assume the Discoverer knows the scheme according to which these paths are picked (but not in which order the nodes appear on a specific path). Thus, if the Discoverer knows the label of an edge $e$, it may deduce which path it belongs to (by calculating $\lfloor\operatorname{\lambda}(e)/n\rfloor$) and what the index of the edge has on the path (by calculating $\operatorname{\lambda}(e)\operatorname{mod}n$).
@@
We only consider the first $(n-1)/2-1$ rounds, since after these, there will have been $n((n-1)/2-1)\in\Omega(n^{2})$ infections. This is where the $1/3$ in the minimum in this theorem comes from, since for large enough $n$, we have $n((n-1)/2-1)>1/3\cdot n^{2}$. This case handled , we assume for the rest of this proof, that during our game there is exactly one infection chain per path.
@@
Intuitively, the Discoverer has information about some of the edges for each path, and the goal of the Discoverer is to find the start of any of the paths. Formally, we prove that the Discoverer must have observed an infection involving the source (i.e., the first node on all paths), that is, it must have watched the first or second node on a path. By construction, an observed infection on one path reveals very limited information about the other paths. An observed infection reveals the positions of the two involved nodes on the current path. This only tells the Discoverer that they may not be adjacent on any other path. Within the first $(n-1)/2-1$ rounds, the Discoverer learns precisely $(n-1)/2-1$ of these pieces of information, thus for each node, there are at least $n/2$ positions on the current path that may be consistent with this information. Also, after $(n-1)/2-1$ rounds, there are at least two nodes on which the Discoverer has no information. Thus, if the Discoverer never observes the first or second node on a path, the adversary may rearrange the paths in such a way, that one of the unobserved nodes becomes the source (it picks whichever node the Discoverer does not claim is the source).

### v2

@@ 1 Introduction
Information, diseases, pollutants—all these things spread through the networks that define how everything from people and servers to our streets and sewers are connected. The field of spreading processes has received substantial attention from researchers as diverse as physicians, physicists, sociologists, and computer scientists. One of the central questions that naturally arises is to find the source of some spreading process given only limited information (e.g., from a small set of disease infection times or from a number of sensors in a sewer network) sir-source-detection; shah2011rumors; sensor-placement; paluch_fast_2018. The current mainstream models in this area, such as the independent cascade model independent-cascade, or the susceptible-infected-resistant (SIR) model Hethcote1989; a-contribution; shah2011rumors incorporate the inherent temporal behavior of the spreading process, that is, infections themselves take time such that a node first gets infected and only then is able to infect other nodes. However, they fail to capture temporal dependencies in the underlying network itself. In many real-world networks the connections change over time (e.g., social networks in which diseases or information spread are highly temporal as the link between two people is only able to transmit information or diseases at specific points in time) and these changes heavily impact the behavior of spreading processes. For this reason, the study of infections in temporal graphs has been initiated. In 2015, Gayraud et al. gayraud-evolving-social-networks, generalized the independent model from static to temporal graphs. Since then, other papers have studied problems related to spreading processes in temporal networks influencers, but, to the best of our knowledge, no one has conducted research on the problem of source detection (SD). We provide the first rigorous definition of SD in temporal graphs and give both algorithms and lower bounds for this problem in a variety of settings. [^1] This work is based on the first author’s master thesis, which is available at https://arxiv.org/abs/2503.13567.
@@ 1.2 Related Work
The SD problem has a rich literature and has even been studied on the SIR model we employ. In 2011, shah2011rumors study the SD problem under the SIR model. In 2016, sir-source-detection extend their work. Other notable algorithmic contributions include the PVTA algorithm proposed by pinto2012locating, which focuses on estimating the source of an infection from observations on a sparse set of nodes. This model has also recently been extended to SD pmlr-v216-berenbrink23a. In 2011, shah2011rumors study the SD problem under the SI model. sir-source-detection extend this work to the SIR model. gayraud-evolving-social-networks, extend the study of influence maximization on both the independent cascade and linear threshold models to temporal graphs (which were popularized by kempe2000connectivity). influencers study a number of variations of the influence maximization problem on temporal graphs under the SIS model (which is closely related to SIR).
@@ 2 Preliminaries
We write $2\operatorname{\mathbb{N}}$ for the set of even natural numbers and $2\operatorname{\mathbb{N}}+1$ for the set of odd natural numbers. Two real-valued functions $f,g\colon\operatorname{\mathbb{N}}\to\operatorname{\mathbb{R}}$ are asymptotically equivalent if $(f(n)/g(n))\xrightarrow{n\to\infty}1$. We use the standard definition of treewidth as presented in cygan2015parameterized.
@@ 4 Randomized Source Detection
Randomizing the behavior of the Discoverer enables us to evade the worst cases and yield good results with high probability. In fact, in this section, we even show that there is a randomized Discoverer that wins the SD game with consistent source behavior on unknown static graphs within $O(n\sqrt{n})$ infections and with constant success probability.
@@
In each round, the strategy picks a node to watch uniformly at random. It terminates when it has observed a node which gets infected while being watched. Let $T$ be the random variable that takes the number of the round in which this happens. Let $a_{1},\dots$ be the nodes infected in the respective rounds. Then the expected number of tolerated infections is $\mathbb{E}\left[\sum_{i=1}^{T}a_{i}\right]=\sum_{i=1}^{\infty}a_{i}\cdot\mathbb{P}\left[i\leq T\right]=\sum_{i=1}^{\infty}a_{i}\cdot\prod_{j=1}^{i-1}\left(\frac{n-a_{j}}{n}\right)$, where the equalities follow from the alternative definition of the expectation and since we require at least $i$ rounds iff the first i1i-1, rounds are unsuccessful. Now, substitute $p_{i}=a_{i}/n$ for all $i$. Then $p_{i}$ is the probability of finishing in round $i$ assuming it is reached. We may also pull $n$ outside the sum to view the number of tolerated infections in each round as a multiple of $n$. This leaves us with $n\sum_{i=1}^{\infty}p_{i}\cdot\prod_{j=1}^{i-1}(1-p_{j})$. Now group the rounds of our strategy and thus the $p_{i}$ by taking consecutive elements until their sum is at least $1/2$ and then start the next group. Formally, let $\{k_{j}\}_{j\in\operatorname{\mathbb{N}}}$ be the (possibly finite) sequence such that $k_{j}$ is the smallest integer such that $\sum_{i=k_{j}}^{k_{j+1}-1}p_{i}\geq 1/2$. Then, by union bound, the probability of finishing within a given group, assuming we reach it, is at least $1/2$. Also, we have that $\sum_{i=k_{j}}^{k_{j+1}-1}p_{i}\leq 3/2$ since all but the last elements together are less than $1/2$ and the last element is at most 1. This is a bound on the tolerated infections (as a multiple of $n$). Therefore, we can bound the above expectation from above by the geometric process that tolerates $3n/2$ infections to perform the rounds in a given group and has probability $1/2$ of finishing within that block. Thus, we can bound above by
$\leq$~ $n\sum_{j=1}^{\infty}\left(\sum_{i=k_{j}}^{i=k_{j+1}-1}p_{i}\right)\cdot\left(\prod_{\ell=1}^{j-1}\left(1-\sum_{i=k_{\ell}}^{i=k_{\ell+1}-1}p_{i}\right)\right)$
@@
Let $0<\varepsilon<1$ be arbitrarily small. For any algorithm solving the SD problem on trees with consistent source behavior, unknown static graph and watching a single node with success probability $p>0$, there is an instance for every number of nodes $n$ such that the price of detection under that algorithm is at least $(1/2)\cdot\sqrt{\ln\left((1-p/(2+\varepsilon))^{-2}\right)}n\sqrt{n}$.
This also gives us a bound on the trade-off between the success probability and the price of detection. For a more intuitive bound, obverse that by simple analytic tools we can deduce that even a price of detection of $pn\sqrt{n}$ is unachievable for any algorithm with success probability $p<0.8724$. {proofsketch} For $n\in\operatorname{\mathbb{N}}^{+}$, we consider the path on $n$ nodes labeled $\operatorname{\lambda}(i,i+1)\coloneqq i$ for $i\in[n-1]$. We assume that there is an algorithm $A$ that wins the game on these graphs within $c\sqrt{n}$ rounds with probability at least $p$. From $A$ we construct algorithm $B$ with strictly stronger capabilities: in each round, $B$ watches the same nodes as $A$ and also all nodes that $A$ has observed to become infected in previous rounds. We argue that for $A$ to win the game, $B$ has to win the game, which happens if $A$ guesses a node which is close to the source. We bound the probability of that event to obtain the result.
@@
First, we describe an infinite family of graphs, then assume there in an algorithm winning the game on all of these instances in $o(n\sqrt{n})$ infections with probability at least $p$. We then derive an algorithm that must solve the problem strictly faster and show that, for all small enough $c$ (which we will give later), there are problem instances where this improved algorithm has probability less than $p$ to require less than $c\sqrt{n}$ rounds.
First of all, let $n\in\operatorname{\mathbb{N}}$. Then set $P_{n}$ the path with $n$ nodes and let the time labeling be $\operatorname{\lambda}(i,i+1)=i$ for $i\in[n-1]$. We shuffle the node indices uniformly at random before providing them to the Discoverer so that they do not reveal the graph For simplicity we will refer to the ordered indices in the rest of the proof. Now let $A$ be any algorithm that solves the SD problem on the set $\{P_{n}\}_{n\in\operatorname{\mathbb{N}}}$. w.l.o.g., we may assume that there is a round in which $A$ watches the root node (if there is no such round, we can modify $A$ such that in a single extra round before submitting its answer, $A$ watches the node it will claim is the source).
@@
By assumption, $A$ requires less than $c\sqrt{n}$ rounds (with probability at least $p$ and for large enough $n$). We may also assume $A$ spends at least one round watching the source once it has found it. Thus, in order for $B$ to have picked the source at the end, $A$ must have picked a node with distance at most $c\sqrt{n}$ from the source at some time . In order for that to happen, $A$ must have picked a node about which it had no information and which has distance at most $c\sqrt{n}$ from the source (by counting the rounds needed to trace an infected node to the source).
Since the nodes about which $A$ has no information were shuffled uniformly at random by the adversary, we can assume that $A$ selects them uniformly at random without replacement. Now, since $2+\varepsilon>2$, the probability of picking a node within distance $c\sqrt{n}$ to the source is smaller than
@@
If we now pick $c<(1/2)\cdot\sqrt{\ln\left(\frac{1}{(1-p/(2+\varepsilon))^{2}}\right)}$ and $n$ large enough such that $c\sqrt{n}\geq 1$, we have that this success probability is less than p. This follows since, by choice, $2c^{2}>\ln(1-p/(2+\varepsilon))$ and thus $e^{-2c^{2}}=\left(e^{-2c}\right)^{c}\geq 1-p/(2+\varepsilon)$. As $\left(1-\frac{(2+\varepsilon)c}{\sqrt{n}}\right)^{\sqrt{n}}$ approaches $e^{-(2+\varepsilon)c}$ and $\frac{n-(2+\varepsilon)c\sqrt{n}}{n}=1-\frac{(2+\varepsilon)c}{\sqrt{n}}$, we have the desired result $1-\left(\frac{n-(2+\varepsilon)c\sqrt{n}}{n}\right)^{c\sqrt{n}}<p$ for large enough $n$. This contradicts the assumption that $A$ had success probability at least $p$, proving that no such algorithm may exist.
@@ 5 When Knowing the Static Graph Helps
The main idea of this reduction is to use the algorithm $A$ as a subroutine by reporting $V\choose 2$} as the edge set to it.
@@
Figure 3: The construction from the proof of Section 5. Discoverer for known graphs AA is used on the completed graph on the right to win the game in the setting of unknown graphs.
@@
This result is achieved by Algorithm 2. It crucially depends on the existence of small separators. Specifically, in a static graph $G$ with treewidth $\mathrm{tw}{}$, there is a $1/2$-balanced separator $S$ (i.e., every connected component in GSG-S has size at most $1/2\left|V(G)\right|$) of size at most $\mathrm{tw}+1$ cygan2015parameterized. See Figure 4 for an illustration. This result also holds if the nodes are weighted, and we ask for a separator to ensure each component has at most half of the weight. Therefore, if we assign either 0 or $1$ as weights, we may pick which nodes we want to evenly distribute on the sides of the separator.
@@ 7 Dynamic Infection Behavior
Let $n\in 2\operatorname{\mathbb{N}}+1$. We construct $G$ as follows: take (n1)/2(n-1)/2 edge-disjoint Hamiltonian paths $P_{1},\dots,P_{(n-1)/2}$ on the $K_{n-1}$ (which exist by axiotis_approx), add a node $s$ and connect $s$ to the first node on each of the Hamiltonian paths. From now on, we interpret each $P_{i}$ as starting at $s$. Order the nodes and the Hamiltonian paths uniformly at random. Note, these cover all edges in $\mathcal{G}$. Write $p\colon E(\mathcal{G})\to[n/2]$ be the function that maps an edge to the Hamiltonian path it belongs to and write $q\colon E(\mathcal{G})\to[n]$ for the function that assigns an edge its index on that path. Then let the Adversary act according to $\operatorname{\lambda}\coloneqq e\mapsto p(e)n+q(e),$ and pick $\delta=1$. See Figure 5 for an illustration of this construction. In each round $i$, the Adversary lets the source infect the whole graph via $P_{i\operatorname{mod}(n-1)/2}$. For the sake of this proof, we may assume the Discoverer knows the scheme according to which these paths are picked (but not in which order the nodes appear on a specific path). Thus, if the Discoverer knows the label of an edge $e$, it may deduce which path it belongs to (by calculating $\lfloor\operatorname{\lambda}(e)/n\rfloor$) and what the index of the edge has on the path (by calculating $\operatorname{\lambda}(e)\operatorname{mod}n$).
@@
We only consider the first (n1)/21(n-1)/2-1 rounds, since after these, there will have been $n((n-1)/2-1)\in\Omega(n^{2})$ infections. This is where the $1/3$ in the minimum in this theorem comes from, since for large enough $n$, we have $n((n-1)/2-1)>1/3\cdot n^{2}$. This case handled , we assume for the rest of this proof, that during our game there is exactly one infection chain per path.
@@
Intuitively, the Discoverer has information about some of the edges for each path, and the goal of the Discoverer is to find the start of any of the paths. Formally, we prove that the Discoverer must have observed an infection involving the source (i.e., the first node on all paths), that is, it must have watched the first or second node on a path. By construction, an observed infection on one path reveals very limited information about the other paths. An observed infection reveals the positions of the two involved nodes on the current path. This only tells the Discoverer that they may not be adjacent on any other path. Within the first (n1)/21(n-1)/2-1 rounds, the Discoverer learns precisely (n1)/21(n-1)/2-1 of these pieces of information, thus for each node, there are at least $n/2$ positions on the current path that may be consistent with this information. Also, after (n1)/21(n-1)/2-1 rounds, there are at least two nodes on which the Discoverer has no information. Thus, if the Discoverer never observes the first or second node on a path, the adversary may rearrange the paths in such a way, that one of the unobserved nodes becomes the source (it picks whichever node the Discoverer does not claim is the source).